Circular No. 7868

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1118+480 C. Zurita, J. Casares, and I. G. Martinez-Pais, Instituto de Astrofisica

de Canarias; and A. Piccioni, S. Bernabei, C. Bartolini, and A. Guarnieri, Bologna University and Bologna Astronomical Observatory, report: “New R-band CCD images of the halo blackhole x-ray transient XTE J1118+480 (IAUC 7389, 7390) have been obtained on Feb. 5 with the 2.5-m NOT telescope (Observatorio del Roque de los Muchachos), on Feb. 22–24 with the 1-m OGS telescope (Observatorio del Teide), and on Mar. 19 with the 1.52-m telescope at Loiano (Osservatorio di Bologna). The system has declined 0.28 mag from its previous reported value (Zurita et al. 2002,

http://xxx.lanl.gov/abs/astro-ph/?0202438) to an average magni-

tude R = 18.93 ± 0.01. The current magnitude is consistent with preoutburst estimates (op.cit.), indicating that the transient has probably reached true quiescence. Lightcurves are dominated by a classical ellipsoidal modulation with night-to-night variations that may be caused by residual superhump contribution (IAUC 7617). The ellipsoidal modulation shows a peak-to-peak amplitude of ∼ 0.26 mag (i.e., ∼ 0.08 mag larger than previous determinations; Zurita et al., op.cit.), as expected by the relative drop in accretion disk brightness.”

SUPERNOVAE 2002bs, 2002bv, 2002bx T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report on spectra (range 370–750 nm) obtained by M. Calkins with the Mt. Hopkins 1.5-m telescope (+ FAST spectrograph). A spectrum of SN 2002bs (cf. IAUC 7861), obtained on Apr. 3.32 UT, confirms it to be a type-Ia supernova; adopting the NASA/IPAC Extragalactic Database (NED) recession velocity of 2895 km/s for the host galaxy, the supernova expansion velocity is ∼ 9500 km/s for Si II (rest 635.5 nm). The spectral-feature age of SN 2002bs (Riess et al. 1997, A.J. 114, 722) is 8 ± 2 days after maximum. A spectrum of SN 2002bx (cf. IAUC 7864), obtained on Apr. 3.15, shows it to be a type-II supernova before maximum; the spectrum consists of a very blue continuum with P-Cyg lines of hydrogen and helium. Adopting the NED recession velocity of 2260 km/s for the host galaxy (IC 2461), the expansion velocity derived from the minimum of the Hβ line is 11 600 km/s. A spectrum of SN 2002bv (cf. IAUC 7864), obtained on Apr. 3.13, shows it to be a type-IIn supernova; the spectrum consists of a blue continuum with hydrogen lines consisting of broad (FWZI ∼ 5400 km/s) and narrow (FWHM ∼ 350 km/s) components.

2002 April 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green